OpenStudy (anonymous):

find the sale price. $150 suite at 15% off

OpenStudy (anonymous):

i wanted to tell 85% of 150 =135$

OpenStudy (anonymous):

how did u get that answer?

OpenStudy (anonymous):

it says 15% is off meaning that the price now is only 85%. get it? so now 85% of 150 will get u the sale price rite? so u hav it.

OpenStudy (anonymous):

@brackett: Are you sure?

OpenStudy (anonymous):

go get 'm brackett!

OpenStudy (anonymous):

or you could simply multiply 150 by .15, then subtract that product from 150.

OpenStudy (anonymous):

150*.15=22.5, then 150-22.5=127.5

OpenStudy (anonymous):

yea 150-150*.15=150(.85)=135....

OpenStudy (anonymous):

oops! calculation error man! but logic is dat simple :/

OpenStudy (anonymous):

method is right, answer is wrong though. i think you need a calculator i get \[.85\times 150=127.50\]

OpenStudy (anonymous):

thanks u all